# cryoSPHERE: Single-particle heterogeneous reconstruction from cryo EM

### v1

@@ 5 Discussion
CryoSPHERE presents several advantages compared to other methods for volume and structure reconstruction:
@@
Deforming a base structure into a density map avoids the computationally expensive $N_{\textrm{pix}}^{2}$ evaluation required by a decoder neural network in methods implicitly parameterising the grid, such as Zhong et al. (2021a); Levy et al. (2022). The neural network size is vastly reduced compared to state of the art methods, which explicitly parameterise the density grid of size $N_{\textrm{pix}}^{3}$, see e.g. Punjani & Fleet (2023). Furthermore, direct deformation of a structure directly avoids the need for subsequent fitting into the recovered density map.
@@
Learning one rigid transformation per segment, where the number of segments is much smaller than the number of residues reduces the dimensionality of the problem. This results in a smaller neural network size compared to approaches acting on each residues, such as Rosenbaum et al. (2021). Rigidly moving large portions of the protein corresponds to low-frequency movements, less prone to noise pollution than the high-frequency movements associated with moving each residue independently. In addition, since our goal is to learn one rotation and one translation per segment, a latent variable of dimension $6\times N_{\textrm{segm}}$ is, in principle, a sufficiently flexible choice to model any transformation of the base structure. Choosing the latent dimension is more difficult for volume reconstruction methods such as (Zhong et al., 2021a).
@@
Section 4 demonstrates cryoSPHERE’s capability to recover continuous and discrete heterogeneity while performing structure reconstruction. It also demonstrate that even though the choice of $N_{\textrm{segm}}$ impact the FSC to the ground truth, cryoSPHERE is able to recover the correct motion for the entire range of $N_{\textrm{segm}}$ values. It is also able to keep the minimum necessary number of domains when the user sets an unnecessarily high $N_{\textrm{segm}}$, as seen in Section 4.1.
While using a base structure $S_{0}$ provides detailed information missed by cryoDRGN, especially evidenced by our better $0.143$ cutoffs, it can introduce bias on medium scale elements, especially if $N_{\textrm{segm}}$ is too small. We alleviate this problem by fitting $S_{0}$ into a volume reconstructed using a simple backprojection algorithm. This, combined with cryoSPHERE, achieves better $0.5$ cutoffs than cryoDRGN, showcasing its effectiveness in debiasing the results. If such a volume is unavailable, simply increasing $N_{\textrm{segm}}$ can reduce the bias.
Our study opens up for significant advancements in predicting protein ensembles and dynamics, critically important for unraveling the complexity of biological systems. By predicting all-atom structures from cryo EM datasets through more realistic deformations, our work lays the foundation for extracting direct insights into thermodynamic and kinetic properties. In the future, we anticipate the ability to predict rare and high-energy intermediate states, along with their kinetics, a feat beyond the reach of conventional methods such as molecular dynamics simulations.
It would be interesting to assess how much our segmentation correlates with bottom-up segmentation into domains conducted on the “omics” scale, see e.g (Lau et al., 2023). To achieve this quantitatively, we would need many examples of moving segments from cryo EM investigations to match the millions of segments from the “omics” studies. Therefore, we leave this investigation to later work.
Finally, this work is an important milestone in showing that one can learn a segmentation of the protein that is intimately linked to the change of conformation of the underlying protein, in an end-to-end fashion. This is in contrast to other works that perform the segmentation based on residue distances of the underlying atomic model only, usually not in an end-to-end way. The application of cryoSPHERE to real dataset requires a number of careful choices that we leave to future work.
@@ Acknowledgement
G.D is financially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) and the Data-Driven Life Science (DDLS) program through the WASP-DDLS collaboration.
This project would not have been possible without the computing resources provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre (Berzelius resource).
We also thank Claudio Mirabello at the National Bioinformatics Infrastructure Sweden at SciLifeLab for providing us access to his AlphaFold installation on Berzelius.
Appendix A Experiments
In this appendix, we provide more details on the experiments of Section 4.
In all the experiments of Section 4, each posed structure is converted into a volume, which is then projected into 2D images according to our image formation model in (2), with $\sigma=1$. The same CTF corruption is applied with parameters specified in Table 1 and finally noise is added to achieve a SNR equal to 0.1. Here, SNR is defined as the ratio of the variance of the images to the variance of the noise.
@@ Appendix A Experiments
A.1 Toy dataset
For this experiment, we predict the phytochrome structure using AlphaFold multimer(Evans et al., 2021) on its amino acid sequence with the UniProt (The UniProt Consortium, 2021) entry Q9RZA4. This protein forms a dimer with 755 residues on each chain. We define two domains for simulation purposes. The first domain comprises the first chain and the first 598 residues of the second chain. The second domain consists of the remaining 157 residues of the second chain. In Figure 3, we present the base structure, the decomposition into the two domains as well as the structures corresponding to the deformed base. These deformations represent the mean rotation of each mode. We rotate the second domain around the $(0,1,0)$ axis, sampling $10^{4}$ rotation angles from the Gaussian mixture:
@@ A.1 Toy dataset
This gives $10^{4}$ structures. For each image, we uniformly sample 15 rotation poses and 15 translation poses on $[-10,10]^{2}$. The structure undergoes rotation, translation, and is then turned into an image according to image formation model 2. Subsequently, the images undergo CTF corruption, and noise is added to achieve $\mathrm{SNR}\approx 0.1$. This process generates a total of $150$k images, each of size $N_{\textrm{pix}}=220$.
We run cryoSPHERE with $N_{\textrm{segm}}=4$ for 48 hours on a single NVIDIA A100 GPU, equivalent to 779 epochs.
Due to computational constraints, the plots in Subsection 4.1 are based on only 10000 images, one per conformation.
Figure 7 illustrates the predicted angles against the latent means, demonstrating that the model effectively learns rotational motion.
Additionally, in Figure 8, we plot the distribution of predicted translation norms for the two segments, the predicted angles of rotation for segment 2 and the distribution of the dot product between the predicted axis of rotation for segment 4 and the true axis of rotation $(0,1,0)$. Our model predicts no translation for each segment and no rotation for the second segment. The axis of rotation is almost exactly recovered.
@@
Figure 8: Toy dataset. Leftmost: Boxplot of the norm of the translations predicted for the fourth segment. Middle left: Boxplot of the norm of the translations predicted for the second segment. Middle right: Boxplot of the predicted angle of rotation for the second segment, in radians. Rightmost: Boxplot of the dot product between the predicted axis of rotation for the fourth segment and the true angle of rotation. CryoSPHERE recovers the right axis of rotation almost perfectly.
A.2 Molecular dynamics dataset
We take the structure of a phytochrome with PDB ID 4Q0J (Burgie et al., 2014) and define two domains: residues 321 to 502 of the first chain and residues 321 to 502 of the second chain. To simulate the dissociation process of the two upper domains, we perform Metadynamics(Barducci et al., 2011) simulations in GROMACS (Abraham et al., 2015; Pronk et al., 2013) with the PLUMED 2 implementation (Tribello et al., 2014). The collective variable chosen is the distance between the self-defined centers of mass (COMs) of the upper domains (residues 321-502 of chain A and B). A 100 ns simulation is conducted using the NpT ensemble, maintaining pressure control through the Parrinello-Rahman barostat. Gaussian deposition occurs every 5000 steps, featuring a height of 0.1 kJ/mol and a width of 0.05 nm. Afterwards, we extract $10^{4}$ structures along the dry trajectory. See Figure 9 for examples of structures. From this, two datasets are generated.
@@ A.2 Molecular dynamics dataset
The first dataset consists $15$ rotation poses uniformly sampled per structure. Translation poses are neglected in this data set, resulting in a total of $150$k images.
The same process is followed to generate the second dataset, with the exception that only one rotation pose per structure is assumed. This results in $10$k images.
@@
For both datasets, we use $N_{\textrm{pix}}=190$, and each pixel is of size $1$Å. The encoder is a 4-hidden-layer neural network with fully connected hidden layers of dimension of $2048,1024,512,512$. The decoder is a 2-hidden-layer neural network with fully connected hidden layers of dimension $350,350$. We set the batch size to $100$ and use the ADAM optimizer (Kingma & Ba, 2017) with a learning rate of $0.0003$, unless otherwise stated. The latent dimension is set to $40$ with $N_{\textrm{segm}}=6$ for all runs, unless stated otherwise. For the 150k dataset, we train for 24 hours on a single NVIDIA A100 GPU, corresponding to 642 epochs. The 10k dataset is trained for 12 hours on the same single GPU, corresponding to 4114 epochs.
@@
To compare cryoSPHERE to cryoDRGN, we train cryoDRGN using default settings - a batch size of 8 and a latent dimension equal to 8 - for the same duration, on the same single GPU used for our algorithm. To maintain consistency when comparing volumes generated from cryoDRGN, our methods, and the ground truth, we convert the structures (both ground truth and predicted with cryoSPHERE) into volumes using the same image formation model employed to generate the dataset.
@@
For computational efficiency, the FSC plots and distances plots are not based on all structures. In the case of the 150k dataset, only one image per structure is used to compute the distances. All images in the 10k dataset are utilized. Consequently, the distance plots are based on 10k images in both cases. For the computation of the FSC curves, we select only 1000 images evenly distributed among these 10000 structures.
A.2.1 150k dataset
Now, we test the robustness of our method to the choice of a base structure.
First, we perform two runs with $N_{\textrm{segm}}=6$. For the first run, we use the same AlphaFold structure as in section 4.3, before it was fitted into the volume obtained by backprojection. In the second run, we use the same AlphaFold structure but fitted as described in Section 4.3.
Figure 10 displays the results for the unfitted run. Despite cryoSPHERE showing excellent agreement with the ground truth in terms of distances compared to cryoDRGN, it misses some medium scale elements, while maintaining a better level of small scale details. We observe underperformance at the $0.5$ cutoff but outperformance at the $0.143$ cutoff. For example, considering the ground truth volumes of the first MD structure and the corresponding reconstructed volumes from cryoSPHERE and cryoDRGN in Figure 13, we note that when only using an AlphaFold structure without postprocessing, our method can be slightly off on some medium-scale elements, especially at the bottom of the protein. As hypothesizing $N_{\textrm{segm}}=6$ is not sufficient to break these elements in specific segments, cryoSPHERE does not move them, retaining the features of the AlphaFold structure, potentially misplaced. In contrast, cryoDRGN seems to be closer to the ground truth on these elements. However, upon closer inspection of the reconstructed volumes, we observe that cryoSPHERE gives much finer details throughout the volume, a charecteristic not replicated by cryoDRGN.
Figure 11 showcases the results for the fitted run. CryoSPHERE now outperforms cryoDRGN at both the $0.5$ and $0.143$ cutoffs while maintaining an excellent agreement with the ground truth. The fitting of the AlphaFold structure has a significantly positive impact on the performances of our method.
For both runs, we compare the reconstructed volumes of the same image from our method to the ground truth in Figure 14. We observe that throughout the volume, the parts that were off compared to the ground truth with the unfitted AlphaFold structure are much better placed with the fitted AlphaFold structure.
Figure 10: MD 150k dataset unfitted AlphaFold structure. Left: Histograms of the distances between the two upper domains. The true distances are in green. The recovered distances are in blue. Middle: Predicted against true distances in Ångström. The black line corresponds to $x=y$. Right: Plot of the Fourier shell correlation curves. The mean across structures for cryoSPHERE is in plain blue. Dotted blue represents the mean $\pm$ one standard deviation. The red, plain line represents the mean of cryoDRGN across volumes, with the dotted red line representing the mean $\pm$ one standard deviation.
Figure 11: MD 150k dataset fitted AlphaFold structure. Left: Histograms of the distances between the two upper domains. The true distances are in green. The recovered distances are in blue. Middle: predicted against true distances in Ångström. The black line corresponds to $x=y$. Right: Plot of the Fourier shell correlation curves. The mean across structures is in plain blue. Dotted blue represents the mean $\pm$ one standard deviation. The red, plain line represents the mean of cryoDRGN across volumes, with the dotted red line representing the mean $\pm$ one standard deviation.
Next, we conduct two additional runs with the 150k dataset: one using the first structure from the MD simulation and the other using the last structure of the MD simulation. Figure 16 displays the results for the first MD structure as base structure, showcasing an excellent agreement with the ground truth. In addition, the FSC curves experience a drop compared to the fitted AlphaFold structure, but they are similar to the unfitted AlphaFold structure. This is likely because using a base structure belonging to the dataset introduces less bias than using an AlphaFold-generated structure with no fitting. Figure 17 presents the same plot, starting from the last MD structures. The conclusions are the same, although we tend to overestimate the distances to some extent. However, the FSC curves are significantly improved and now match the fitted AlphaFold structure run.
Figure 18 shows the segment decompositions recovered by cryoSPHERE for all these runs, revealing their overall similarity. Notably, two small segments are identified in the helices linking the upper part to the lower one. This is not problematic since our decomposition is soft, allowing for smooth transitions between no motion and the motion of the upper domains.
As expected, the quality of the results depends on the quality of the base structure used. Fortunately, AlphaFold can provide reasonable structures that we can further fit into density maps obtained with traditional methods. This approach combines the strength of both world: the more static parts of a protein can exhibit good resolution with traditional tools, and we can fit in the corresponding part of an AlphaFold structure, thereby debiasing the method to some extent. On the more mobile parts where traditional methods may have lower resolution, AlphaFold can provide improved resolution.
In a final experiment, we attempt to increase the number of segments to $N_{\textrm{segm}}=16$ using the unfitted AlphaFold structure as the base structure. We run our algorithm for 24 hours on the same single GPU used in previous experiments. Figure 12 demonstrates an excellent agreement with the ground truth in terms of distances. The FSC curves indicate that increasing the number of segments actually improves the results, approaching the performance of the run with the fitted AlphaFold structure, despite using the unfitted AlphaFold structure. Figure 15 visually confirms this improvement. As expected, more segments allow the model to adjust better to the data, reducing the bias introduced by the AlphaFold structure. However, this improvement comes at a cost: in 24 hours, cryoSPHERE with the unfitted AlphaFold structure and 6 segments makes 642 epochs, whereas the run with 16 segments completes 421 epochs.
Figure 12: 150k dataset with the unfitted AlphaFold structure and 16 segments. Left: Histograms of the distances of the two upper domains. The true distances are in green. The recovered distances are in blue. Middle: predicted against true distances in Ångström. The black line corresponds to $x=y$. Right: Fourier shell correlation curves for cryoSPHERE. The mean of of the unfitted AlphaFold with 6 segments is in blue. Dotted blue represents the mean $\pm$ one standard deviation. The red, plain line represents the mean of the unfitted AlphaFold run with 16 segments, with the dotted red line representing the mean $\pm$ one standard deviation. The green, plain line represents the mean of the run with the fitted AlphaFold structure and 6 segments.

### v2

@@ 6 Discussion
CryoSPHERE presents several advantages compared to other methods for volume and structure reconstruction.
@@
Deforming a base structure into a density map avoids the computationally expensive $N_{\textrm{pix}}^{2}$ evaluation required by a decoder neural network in methods implicitly parameterising the grid, such as Zhong et al. (2021a); Levy et al. (2022). Furthermore, direct deformation of a structure directly avoids the need for subsequent fitting into the recovered density map.
@@
Learning one rigid transformation per segment, where the number of segments is much smaller than the number of residues, reduces the dimensionality of the problem. This results in a smaller neural network size compared to approaches acting on each residues, such as Rosenbaum et al. (2021). Rigidly moving large portions of the protein corresponds to low-frequency movements, less prone to noise pollution than the high-frequency movements associated with moving each residue independently. In addition, since our goal is to learn one rotation and one translation per segment, a latent variable of dimension $6\times N_{\textrm{segm}}$ is, in principle, a sufficiently flexible choice to model any transformation of the base structure. Choosing the latent dimension is more difficult for volume reconstruction methods such as (Zhong et al., 2021a).
@@
Section 5 and Appendix B.2 demonstrate cryoSPHERE’s capability to recover conformational heterogeneity while performing structure reconstruction. The division into $N_{\textrm{segm}}$ is learned from the data and only marginally impacts the FSC to the ground truth. Moreover, cryoSPHERE recovers the correct motion for the entire range of $N_{\textrm{segm}}$ values and is able to keep the minimum necessary number of domains when the user sets it too high (Appendix B.1).
Structural restraints allow interpretation of low SNR datasets:
It is evident that structural restraints as implemented in cryoSPHERE (this work) and cryoSTAR provide additional restraints that pure volume methods (i.e. cryoDRGN) lack, thus giving better reconstructions for high noise data sets. The additional restraints may introduce bias, which needs to be alleviated using a backprojection algorithm. This, combined with cryoSPHERE’s latent space, achieves better $0.5$ cutoffs than cryoDRGN, indicating its effectiveness in resolving conformational heterogeneity and debiasing the results. If such a volume is unavailable, simply increasing $N_{\textrm{segm}}$ can reduce the bias. As a note of caution we find that for most dynamic protein studies here (the Pfr state of the phytochrome), we find that volume-based debiasing fails because of the very low electron density levels in the reconstructions. Here, other metrics should be developed in the future.
Summary:
Our study opens up for significant advancements in predicting protein ensembles and dynamics, critically important for unraveling the complexity of biological systems. By predicting all-atom structures from cryo-EM datasets through more realistic deformations, our work lays the foundation for extracting direct insights into thermodynamic and kinetic properties. This work is an important milestone in showing that one can learn a segmentation of the protein that is intimately linked to the change of conformation of the underlying protein, in an end-to-end fashion. In the future, we anticipate the ability to predict rare and high-energy intermediate states, along with their kinetics, a feat beyond the reach of conventional methods such as molecular dynamics simulations.
It would be interesting to assess how much our segmentation correlates with bottom-up segmentation into domains conducted on the “omics” scale, see e.g. Lau et al. (2023). To achieve this quantitatively, we would need many examples of moving segments from cryo-EM investigations to match the millions of segments from the “omics” studies. Therefore, we leave this investigation to later work.
@@ Acknowledgement
This work was financially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) and the Data-Driven Life Science Program (DDLS) funded by the Knut and Alice Wallenberg Foundation through the WASP-DDLS collaboration, the Swedish Research Council (project no: 2020-04122, 2024-05011), and the Excellence Center at Linköping–Lund in Information Technology (ELLIIT). Our computations were enabled by the Berzelius resource at the National Supercomputer Centre, provided by the Knut and Alice Wallenberg Foundation.
We thank Claudio Mirabello at the National Bioinformatics Infrastructure Sweden at SciLifeLab for providing us access to his AlphaFold installation on Berzelius and Nancy Pomarici for providing input files and explanation of the metadynamics simulation.
7 Reproducibility statement
As part of the current paper, we provide a github link to the source code in Section 1. We also describe in detail how we generate the synthetic datasets in Appendix B.2 and the hyperparameters chosen to run cryoStar, cryoSPHERE and cryoDRGN in Appendix B for each of the experiments.
Appendix A Method
In this appendix we provide more details on cryoSPHERE.
A.1 Loss
In our experience it is not necessary to add any regularization term to the loss Equation 4, except for the datasets featuring a very high level of conformational heterogeneity. In that case, we offer the option to add a continuity loss, which prevents the network from breaking the protein, as well as a clashing loss, which prevents clashing between different residues. The idea of a continuity loss has been introduced in (Jumper et al., 2021) and exploited in cryoEM in (Li et al., 2023; Rosenbaum et al., 2021). The clashing loss has also been introduced in (Jumper et al., 2021) and has been exploited in (Li et al., 2023).
For two subsequent residues belonging to a same chain, we define the continuity loss as:
$\mathcal{L}_{\mathbf{cont}}=\dfrac{1}{N_{\mathbf{cont}}}\sum_{i}^{N_{\mathbf{cont}}}||d_{i}-\hat{d}_{i}||^{2}$ (9)
where $N_{\mathbf{cont}}$ is the number of pairs of residues that are subsequent in the entire protein, $d_{i}$ is the distance between the two residues in pair $i$ in the base structure $\mathcal{S}_{0}$ and $\hat{d}_{i}$ is the predicted distance for the corresponding pair. The form of our continuity loss is similar to (Li et al., 2023).
We define the clashing loss as:
$\mathcal{L}_{\mathbf{clash}}=\dfrac{1}{N_{\mathbf{clash}}}\sum_{i}^{N_{\mathbf{clash}}}||\hat{d}_{i}-k_{\mathrm{clash}}||^{2}$ (10)
where $N_{\mathbf{clash}}$ denotes the number of clashing residues in the protein, where two residues are said to clash if $\hat{d}_{i}<k_{\mathrm{clash}}$, with $k_{\mathrm{clash}}=4$ by default. For very large proteins with a high number of residues, computing this clashing loss is impractical. In that case, we compute it for pairs of residues that are distant from $4$ to $10$Å in the base structure $S_{0}$. Note that our clashing loss takes into account all of the residues. This is not the case of Li et al. (2023), who implements a similar form as our clashing loss for big proteins while they describe the same loss as we do their paper.
Appendix B Experiments
In this appendix, we provide more details on the experiments of Section 5. We followed the same approach to create all the images of the synthetic datasets. We first pose the ground truth structure, which we then convert into a volume, which we then project into a 2D image according to our image formation model in (2) with $\sigma=2$. After that, we corrupt all the images according to the same CTF parameters described in Table 1. Finally Gaussian noise is added to achieve different SNRs. Here, SNR is defined as the ratio of the variance of the images to the variance of the noise. In this context, the poses are assumed to be exactly known. However, since we use a structure $S_{0}$, which is different from the structures used to generate the datasets (unless stated otherwise), these poses can only be an approximation. This will not be the case of cryoDRGN, for which these poses will indeed be exact, as this method does not use a base structure. Consequently, in this context, the comparison may introduce a bias in favor of cryoDRGN.
@@ Appendix B Experiments
B.1 Toy dataset
For this experiment, we predict the phytochrome structure using AlphaFold multimer Evans et al. (2021) on its amino acid sequence with the UniProt The UniProt Consortium (2021) entry Q9RZA4. This protein forms a dimer with 755 residues on each chain. We define two domains for simulation purposes. The first domain comprises the first chain and the first 598 residues of the second chain. The second domain consists of the remaining 157 residues of the second chain. We rotate the second domain around the $(0,1,0)$ axis, sampling $10^{4}$ rotation angles from the Gaussian mixture:
@@ B.1 Toy dataset
In Figure 8, we present the base structure, the decomposition into the two domains as well as the structures corresponding to the deformed base. These deformations represent the mean rotation of each mode. This gives $10^{4}$ structures. For each structure, we uniformly sample 15 rotation poses and 15 translation poses on $[-10,10]^{2}$. The structure undergoes rotation, translation, and is then turned into an image according to image formation model 2. Subsequently, the images undergo CTF corruption, and noise is added to achieve $\mathrm{SNR}\approx 0.1$. This process generates a total of $150$k images, each of size $N_{\textrm{pix}}=220$.
We run cryoSPHERE with $N_{\textrm{segm}}=4$ for 48 hours on a single NVIDIA A100 GPU, equivalent to 779 epochs. The encoder has 4 hidden layers of size 2048, 1024, 512, 512 and the decoder has two hidden layers of size 350, 350.We use a learning rate of $0.00003$ for the parameters of the decoder and encoder and a learning rate of $0.0003$ for the segmentation GMM parameters.
Due to computational constraints, the plots in this section are based on only 10000 images, one per conformation.
Moreover, Figure 8 shows that most of the predicted angles of rotation for the fourth segment are in excellent agreement with the ground truth structural changes. In addition, the predicted translations for both segments are close to 0, the predicted axis of rotation of the moving segments is close to $(0,1,0)$ and the predicted rotation angles for non moving segment are null, see Figure 10.
Finally, Figure 9 illustrates the predicted angles against the latent means, demonstrating that the model effectively learns rotational motion.
@@
Figure 10: Toy dataset. Leftmost: Boxplot of the norm of the translations predicted for the fourth segment. Middle left: Boxplot of the norm of the translations predicted for the second segment. Middle right: Boxplot of the predicted angle of rotation for the second segment, in radians. Rightmost: Boxplot of the dot product between the predicted axis of rotation for the fourth segment and the true axis of rotation. CryoSPHERE recovers the right axis of rotation almost perfectly.
B.2 Molecular dynamics dataset
We take the structure of a phytochrome with PDB ID 4Q0J Burgie et al. (2014) and define two domains: residues 321 to 502 of the first chain and residues 321 to 502 of the second chain. To simulate the dissociation process of the two upper domains, we perform MetadynamicsBarducci et al. (2011) simulations in GROMACS Abraham et al. (2015); Pronk et al. (2013) with the PLUMED 2 implementation Tribello et al. (2014). The collective variable chosen is the distance between the self-defined centers of mass (COMs) of the upper domains (residues 321-502 of chain A and B). A 100 ns simulation is conducted using the NpT ensemble, maintaining pressure control through the Parrinello-Rahman barostat. Gaussian deposition occurs every 5000 steps, featuring a height of 0.1 kJ/mol and a width of 0.05 nm. Afterwards, we extract $10^{4}$ structures along the dry trajectory. See Figure 11 for examples of structures. The closed conformation is the starting conformation of the MD simulation and the most open one corresponds to the end.
@@ B.2 Molecular dynamics dataset
For each structure, we sample $15$ rotation poses uniformly together with translations uniformly on $[-10,10]^{2}$. This results in a dataset of $150$k images. We use $N_{\textrm{pix}}=190$, and each pixel is of size $1$Å. We finally add Gaussian noise to create three datasets: a $\mathrm{SNR}=0.1$, a $\mathrm{SNR}=0.01$ and $\mathrm{SNR}=0.001$ dataset. The results for $\mathrm{SNR}=0.001$ are described in Section 5.1 and this appendix describes the results for the $\mathrm{SNR}=0.1$ and $\mathrm{SNR}=0.01$ datasets.
@@
For cryoSPHERE on all three datasets, the encoder is a 4-hidden-layer neural network with fully connected hidden layers of dimension of $512,256,64,64$. The decoder is a 2-hidden-layer neural network with fully connected hidden layers of dimension $512,512$. We set the batch size to $128$ and use the ADAM optimizer Kingma & Ba (2017) with a learning rate of $0.00003$ for the encoder and decoder parameters, while we set a learning rate of $0.0003$ for the GMM segmentation parameters. The latent dimension is set to $8$ and we run the program with $N_{\textrm{segm}}=10$, $N_{\textrm{segm}}=20$ and $N_{\textrm{segm}}=25$ for all datasets. We train for 24 hours on a single NVIDIA A100 GPU.
@@
We use the default parameters for both cryoStar and cryoDRGN but disable the structural loss of cryoStar, except for the elastic network loss. We train both methods for 24 hours on the same GPU as cryoSPHERE. To maintain consistency when comparing volumes generated from cryoDRGN, our methods, cryoStar and the ground truth, we convert the structures (ground truth, predicted with cryoStar and predicted with cryoSPHERE) into volumes using the same image formation model employed to generate the dataset, see 1. Since cryoStar also proposes a volume method similar to cryoDRGN, after training the structure method of cryoStar for 24 hours, we train their volume methods for the same amount of time. We report the results for this volume method in this appendix.
Since cryoSPHERE and cryoStar are structural methods, we can compute the predicted distance between the two domains defined earlier for each image, both for cryoStar and cryoSPHERE.
@@
For computational efficiency, the FSC plots and distances plots are not based on all structures: only one image per structure is used to compute the distances. Consequently, the distance plots are based on 10k images. For the computation of the FSC curves, we select only 1000 images evenly distributed among these 10000 structures.
B.2.1 SNR $0.1$
This subsection describes the results for cryoSPHERE, cryoStar and cryoDRGN on our molecular dynamics simulation dataset with SNR $0.1$. We use $N_{\textrm{segm}}=25$ in this section.
Figures 13 and 14 show that both cryoSPHERE and cryoStar recover the ground truth distribution of distances very well. They are also able to identify the correct conformation conditionally on an image, as illustrated by the predicted versus true distances plot.
Figure 15, shows that both cryoStar and cryoSPHERE outperform cryoDRGN at both the $0.5$ and $0.143$ cutoffs. It seems cryoStar slightly outperforms cryoSPHERE at both cutoffs. This might be because the SNR is rather high, hence moving each residue individually offers a greater flexibility than moving segments, while the risk of overfitting is low. In addition, we can see that the volume method of cryoStar perform similarly, if not worse, than cryoDRGN. That seems to indicate that this volume method does not benefit from the information gained by the structural method. Figure 12 shows three examples of volumes predicted by the volume method of cryoStar, together with the corresponding ground truth.
Figure 16 and 17 show a set of predicted structures compared to the ground truth for cryoSPHERE and cryoStar. Both methods are able to recover the ground truth almost perfectly.
Figure 18 shows examples of cryoDRGN predicted volume together with the corresponding ground truth. The method is able to recover the ground truth volumes almost perfectly.
Figure 12: SNR $0.1$. Example volumes produced by the volume method of cryoStar with the corresponding ground truth. Top: ground truth. Bottom: corresponding volumes predicted by the volume method of cryoStar.
Figure 13: Results for cryoSPHERE on SNR $0.1$ with $N_{\textrm{segm}}=25$. Left: distribution of distances predicted by cryoSPHERE compared to the ground truth distribution. Middle: true versus predicted distances for cryoSPHERE. Right: segments decomposition.
Figure 14: Results for cryoStar on SNR $0.1$. Left: distribution of distances predicted by cryoStar compared to the ground truth distribution. Right: true versus predicted distances for cryoStar
Figure 15: SNR $0.1$. Mean Fourier shell correlation ($\pm$ std) comparison for cryoSPHERE with $N_{\textrm{segm}}=25$, cryoStar and cryoDRGN. Top left: cryoSPHERE versus cryoDRGN. Top right: cryoSPHERE vs cryoStar. Bottom left: cryoDRGN vs cryoStar. Bottom right: cryoStar volume method vs cryoDRGN
Figure 16: SNR $0.1$. Examples of reconstructed structures by cryoSHERE. Blue is predicted, green is ground truth. CryoSPHERE is able to recover the right conformation. Left to right: image number 11, 5001 and $9$.
Figure 17: SNR $0.1$. Examples of reconstructed structures by cryoStar. Red is predicted, green is ground truth. cryoStar is able to recover the right conformation. Left to right: image number 11, 5001 and $9$.